What is the de Broglie wavelength for a particle moving with

speed 2.0 × 106 m/s if the particle is (a) an electron, (b) a proton,

and (c) a 0.20-kg ball?


1
Expert's answer
2021-05-17T18:57:13-0400

Gives

Particle velocity (v)="2\times10^6m/sec"

(a)

Particle electron

"\lambda_e=\frac{h}{m_ev} \rightarrow(1)"

For electron

"h=6.625\times10^{-34}Jsec"

"m_e=9.1\times10^{-31}kg"

Put value equation (1)

"\lambda=\frac{6.625\times10^{-34}}{9.1\times10^{-31}\times2\times10^6}"

"\lambda= 3.65\times10^{-10}meter=3.65\text{\AA}"

(b)

Particle proton

"\lambda_{p}=\frac{h}{m_pv} \rightarrow(2)"

"\lambda_{p}=\frac{6.625\times10^{-34}}{1.67\times10^{-27}\times2\times10^6}"

"\lambda_{p}=1.983\times10^{-13}meter"

(c)"\lambda_{ball}=\frac{h}{m_{ball}v} \rightarrow(3)"

"m_{ball}=0.20kg"

"\lambda_{ball}=\frac{6.625\times10^{-34}}{0.20\times2\times10^6}"

"\lambda_{ball}=1.65\times10^{-39}meter"